4. The Laffer curve Governments often place so-called sin taxes on goods or services such as cigarettes, alcohol, and pornogrGraph Input Tool Market for Adult DVDs 40 36 32 28 24 20 16 12 Quantity (DVDs) Demand Price (Dollars per DVD) Tax Wedge (DollSuppose the government imposes an 8-per-DVD tax on suppliers At this tax amount, the equilibrium quantity of adult DVDs is DV1600 T 1440 Laffer Curve 1280 1120 2 960 D 800 640 F 480 320 160 0 4 812 16 20 24 28 32 36 40 TAX (Dollars per DVD)Suppose the government is currently imposing a $24-per-DVD tax on adult DVDs. True or False: The government can raise its tax1600 T 1440 Deadweight Loss 1280 1120 960 -800 640 480 320 160 0 4812 16 20 24 28 32 36 40 0 TAX (Dollars per DVD) As the tax
Answer #1

Suppose the government imposes $8 per DVD tax on suppliers.

As a result, the price buyers paid equals to $24 and price sellers receives equals to $16, At this tax amount, the equilibrium quantity of adult DVDs is 80 DVDs and government collects ($8)(80)= $640 in revenue.

To plot the laffer curve, we have to calculate the tax revenue at different tax rates, this is shown in the following table:

Tax Quantity after tax Tax revenue
0 100 0
8 80 640
16 60 960
20 50 1000
24 40 960
32 20 640
40 0 0

By plotting these points, we get the laffer curve as shown below:

-1.602 12o 4ba goa

Suppose the government is currently imposing a $24 per DVD tax on adult DVDs.

FALSE because the government cannot raise its tax revenue by increasing the per unit tax on adult DVDs . As we can see in the above table or graph ,that when tax is $24 per DVD ,then tax revenue =$960 and when tax is $32 ,then tax revenue is $640.

Deadweight loss is the area of the triangle = (0.5)(Reduction in quantity after tax)(Tax amount)

Tax Reduction in quantity Deadweight loss
0 0 0
16 (100-60)=40 (0.5)(40)(16)= 320
32 (100-20)=80 (0.5)(80)(32)=1280

By plotting these points ,we get the deadweight loss curve as shown below:

Yo 33%

As the tax per DVD increases,the deadweight loss increases by greater and greater amount as shown in the figure above.